# Adapting Large Language Models for Document-Level Machine Translation

### v2

@@ Abstract
Large language models (LLMs) have made significant strides in various natural language processing (NLP) tasks. Recent research shows that the moderately-sized LLMs often outperform their larger counterparts after task-specific fine-tuning. In this work, we delve into the process of adapting LLMs to specialize in document-level machine translation (DocMT) for a specific language pair. Firstly, we explore how prompt strategies affect downstream translation performance. Then, we conduct extensive experiments with two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Our findings indicate that in some cases, these specialized models even surpass GPT-4 in translation performance, while they still significantly suffer from the off-target translation issue in others, even if they are exclusively fine-tuned on bilingual parallel documents. Furthermore, we provide an in-depth analysis of these LLMs tailored for DocMT, exploring aspects such as translation errors, discourse phenomena, training strategy, the scaling law of parallel documents, additional evaluation on recent test sets, and zero-shot crosslingual transfer. Our findings not only shed light on the strengths and limitations of LLM-based DocMT models but also provide a foundation for future research.
[…]
@@ 1 Introduction
In this study, we analyze moderately-sized LLMs (with $7B$ parameters) across 18 translation tasks involving nine language pairs. We fine-tune three LLMs using Parameter-Efficient Fine-Tuning (PEFT) and Fully Fine-Tuning (FFT). Comparisons with state-of-the-art translation models, using metrics like $s$BLEU, $d$BLEU, and COMET, confirm the superior translation capabilities of LLMs after fine-tuning. However, we identify a significant issue of off-target translations, observed even after exclusive fine-tuning on bilingual corpora. Additionally, we present an in-depth analysis of our LLM-based DocNMT models from various perspectives: translation error distribution, discourse phenomena, training strategy, the scaling law of parallel documents, additional evaluations on WMT2023 test sets, and zero-shot cross-lingual transfer, aiming to enhance understanding and efficacy of LLMs in DocMT tasks.
@@
  • Selective Excellence in Translation Tasks: Our findings show that our moderately-sized LLMs outperform gpt-4-turbo in certain translation tasks, but struggle in others due to the off-target translation issue. Despite this, our DocMT models exhibit better context awareness and fewer errors, while maintaining comparable performance.
@@ 2 Related Work
Large language models (LLMs) have demonstrated remarkable proficiency across a wide range of Natural Language Processing (NLP) tasks (Brown et al., 2020; Chowdhery et al., 2022; Scao et al., 2022; Anil et al., 2023; Touvron et al., 2023a, b). Furthermore, recent research has shown that supervised fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) can significantly enhance their performance when following general language instructions Weller et al. (2020); Mishra et al. (2022); Wang et al. (2022); Shen et al. (2023); Li et al. (2023); Wu and Aji (2023). More recently, there is a growing body of work exploring the translation capabilities of LLMs (Lu et al., 2023; Zhang et al., 2023; Xu et al., 2023; Robinson et al., 2023). However, it is important to note that these efforts have primarily focused on sentence-level machine translation (SenMT) and have not delved into document-level machine translation (DocMT). A noteworthy study in DocMT is conducted by Wang et al. (2023b), where they investigate the document-level translation capabilities of gpt-3.5-turbo, making it the most closely related work to our work.
@@ 3.1 Two-Stage Training
DocMT approaches typically begin by pre-training the translation model on sentence-level parallel corpora, subsequently refining it through fine-tuning on document-level parallel corpora (Voita et al., 2019; Maruf et al., 2019; Ma et al., 2020; Sun et al., 2022; Wu et al., 2023). More recently, Xu et al. (2023) propose a two-stage training strategy, which initially involves fine-tuning a LLM on monolingual text, followed by a second fine-tuning phase on parallel text. Given that most state-of-the-art open-sourced LLMs are trained on English-centric corpora, our approach begins with the fine-tuning of a LLM on monolingual documents, followed by fine-tuning on parallel documents. Following Xu et al. (2023), we omit the step of fine-tuning on sentence-level parallel datasets.
@@ 3.3 Models
  • State-of-the-art LLMs: For our baseline LLMs in the context of DocMT, we utilize gpt-3.5-turbo and gpt-4-turbo. [^2] Model signatures: gpt-3.5-turbo-1106 and gpt-4-1106-preview. We use the Prompt 4 as detailed in 0(d) during the translation process.
  • Our re-implemented DocMT models: We conduct full fine-tuning on the concatenation-based DocMT model (Tiedemann and Scherrer, 2017), as well as several recent DocMT baselines (Sun et al., 2022; Wu et al., 2023, 2024), initialized with mT5 (Xue et al., 2021). These models are available with parameters of 300M, 580M, and 1.2B, representing the strong DocMT baseline.
@@ 3.4 Evaluation
We use beam search with the beam size of 5 during translation. As shown in 0(d), previous translations serve as the context for the current translation, so the test examples are translated in their original order, beginning with the first sentence free from context.
@@ 4 A Preliminary Study on Prompts
Table 2: Overall performance on IWSLT2017. # of param. indicates the number of parameters of the model. # of train. param. indicates the number of trainable parameters of the model. All the LLM approaches use Prompt 4 (0(d)) during inference. Best results are highlighted in bold. Best results in each group are underlined.
@@ Results
Our investigation analyzes prompt variations using three PEFT models (L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA) on four English-centric translation tasks involving German and Chinese. Overall results are presented in Table 1. Comparing Prompt 1 (0(a)) and Prompt 2 (0(b)), we find that models fine-tuned with Prompt 2 generally outperform those with Prompt 1, indicating Prompt 2’s effectiveness in enhancing LLM performance. Regarding our second research question (0(a) vs. 0(c)), we observe varied performance. L-7B-LoRA and B-7B-LoRA perform better with Prompt 3, while V-7B-LoRA performs better with Prompt 1. These results highlight varying impacts of prompt variations across models and suggest natural language instructions are less effective when using instruction-tuned language models as model backbones. Finally, LLMs with Prompt 4 (0(d)) achieve the best overall performance, suggesting a positive compound effect of context structure and instructions.
@@ Conclusion
As expected, the prompt plays a significant role in LLM performance. A well-structured prompt, which combines an appropriate context structure and natural language instructions, can significantly boost model performance. In this work, we use Prompt 4 (0(d)) in our other experiments, unless otherwise mentioned.
@@ 5 Main Results
In our results presented in Table 2, we observe that gpt-4-turbo and gpt-3.5-turbo significantly outshine all other models in performance. Notably, the NLLB variants, which are trained on vast amount of parallel sentence pairs, also demonstrate superior performance among specialized machine translation (MT) models. In the context of DocMT, conventional DocMT models still outperform our LLM-based DocMT models for translations from English to other languages when evaluated using standard MT metrics. Conversely, for translations from other languages to English, our LLM-based DocMT models perform on par or better than conventional DocMT models in $\mu_{s\textsc{BLEU}}$ and $\mu_{d\textsc{BLEU}}$ metrics, while those conventional DocMT models maintain superior performance in $\mu_{\textsc{COMET}}$.
@@
As indicated in Table 2, our models incorporating LoRA typically outperform fully fine-tuned (FFT) LLMs. However, an exception is observed where V-7B-fft outperforms V-7B-LoRA in translating from other languages to English. This discrepancy is likely attributable to overfitting. In scenarios of extensive fine-tuning with a large corpus of parallel documents, the full fine-tuning of all parameters often leads to rapid overfitting on the training dataset. In contrast, the parameter-efficient fine-tuning approach, exemplified by LoRA, updates only a select number of parameters, effectively preventing the models from overfitting the training set. Furthermore, we observe that the L-7B and V-7B models exhibit comparable performance, suggesting that initializing with instruction-tuned models does not always enhance task-specific performance.
@@
We present the results for the translation tasks from other languages to English in Figure 2. Regarding the readability of the figures, we present only the results provided by our models using LoRA. Our LLM-based DocMT models exhibit superior performance, sometimes even surpassing gpt-4-turbo in certain translation tasks. However, they fail completely in others. A manual review of translation tasks where our LLM-based DocMT models fail reveals that the primary cause of failure is off-target translation. We provide an in-depth analysis of the off-target translation problem in Section 6. A complete breakdown of the results is in Appendix E.
002002002002004004004004006006006006008008008008001,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 8002,00020002{,}0002 , 0002,20022002{,}2002 , 2002,40024002{,}4002 , 400MistranslationOvertranslationUndertranslationAdditionOmissionGrammarCohesionCoherenceInconsistent styleMultiple terms in translation1,39513951{,}3951 , 3957177177177171,22412241{,}2241 , 2248408408408404844844844842342342342347097097097094544544544547257257257253463463463461,35613561{,}3561 , 3567177177177171,35813581{,}3581 , 3588638638638636166166166162662662662667277277277275655655655657717717717713393393393391,71217121{,}7121 , 7127157157157159779779779772452452452454174174174174864864864868048048048045925925925928248248248242732732732732,00220022{,}0022 , 0028368368368361,55115511{,}5511 , 551586586586586868868868868474474474474933933933933717717717717854854854854386386386386# of occurences
Figure 3: Error type analysis given by gpt-4-turbo for translations from English to German, Romanian, and Chinese. The error types in orange are context-dependent. We omit those error types that are rare or almost never occur.
6 Analyses
In this section, we investigate the off-target problem and leverage gpt-4-turbo to analyze the translation errors. We also explore discourse phenomena, the training strategy, and the scaling law of parallel documents. Furthermore, we conduct additional evaluations on recent test sets from WMT2023 and examine crosslingual transfer.
@@ 6 Analyses
Table 3: Off-target rate (%) provided by our LLM-based DocMT models for translation tasks from selective languages to English. $\mu_{\%}$ indicates the average off-target rate across all nine language pairs. A lower off-target rate indicates better performance.
@@
In Figure 2, our LLM-based DocMT models excel in some translation tasks but struggle in others due to off-target translation issues. We investigate this problem using the fasttext library (Bojanowski et al., 2017) to identify translation languages and quantify off-target rates, which represent the proportion of translations that are off-target. Results are presented in Table 3, with off-target rates reaching up to $98.3\%$ in failing tasks. Notably, only B-7B-LoRA consistently maintains low off-target rates, likely due to Bloom-7B’s multilingual pre-training. These findings shed light on the main reason of translation failures in LLM-based DocMT models, offering insights for future research. Detailed off-target rates are provided in Appendix F.
@@
To comprehensively understand the translation capabilities of our LLM-based DocMT models, we select specific error types from the Multidimensional Quality Metrics (MQM) framework (Burchardt, 2013). Kocmi and Federmann (2023) demonstrate gpt-4 is capable of identifying error spans and achieving state-of-the-art MT evaluation accuracy, so we leverage gpt-4-turbo to analyze the translation errors of the text translated by these models. We focus on four models due to resource constraints: L-7B-LoRA, L-7B-fft, Doc2Doc-mT5-1.2B, and GoogleTrans, assessing translations from English to German, Romanian, and Chinese. The error identification prompt is detailed in Appendix D, and we present the frequency of error types in Figure 3. Notably, most errors are limited to individual sentences. Despite similar scores in metrics such as $s$BLEU, $d$BLEU, and COMET among the models, our LLM-based DocMT models (L-7B-LoRA and L-7B-fft) exhibit fewer context-independent and context-dependent errors. This highlights a limitation in current evaluation metrics, suggesting they may not sufficiently assess document-level translations. It also indicates that fine-tuning LLMs for machine translation holds promise for enhancing DocMT performance.
Acc. er es sie
Doc2Doc-mT5-1.2B 77.0 68.7 89.0 73.5
MR-Doc2Sen-mT5 59.9 48.9 91.4 39.4
MR-Doc2Doc-mT5 78.2 67.5 91.1 76.1
DocFlat-mT5 78.0 68.9 90.1 75.1
IADA-mT5 79.1 70.0 89.8 77.6
L-7B-LoRA 83.1 77.2 96.6 75.4
L-7B-fft 81.1 70.2 96.9 76.2
B-7B-LoRA 75.5 56.2 95.1 75.1
B-7B-fft 68.3 50.8 95.5 58.5
V-7B-LoRA 84.9 78.4 96.2 80.1
V-7B-fft 84.4 76.3 96.4 80.5
Table 4: Accuracy (in %) on the English-German contrastive test set. Best results are highlighted in bold.
@@
To evaluate our LLM-based DocMT model’s ability to leverage contextual information, we assessed it using the English-German contrastive test set by Müller et al. (2018). This evaluation tests the model’s accuracy in selecting the correct German pronoun (“er”, “es”, and “sie”) from multiple translation options. Results, shown in Table 4, reveal that models initialized with Llama2-7B and Vicuna-7B outperform Doc2Doc-mT5-1.2B, while Bloom-7B-initialized models perform worse, indicating that contextual understanding is mostly acquired during pre-training, as detailed by Scao et al. (2022) due to the lack of German text in Bloom pre-training.
@@
Table 5: Comparison between two-stage and three-stage training strategies. The results of the two-stage strategy are given by L-7B-fft. For the three-stage training strategy, we fine-tune all the model parameters of Llama2-7B in all three stages.
1102030405060708090100404040406060606080808080Percentage (%) of training dataCOMET
@@
$\mu_{\Delta}$ Ar De Fr It Ja Ko Nl Ro Zh
L-7B-LoRA +29.4 +36.3 +38.8 +37.2 +32.1 +15.9 +17.1 +21.7 +35.8 +29.5
L-7B-fft +29.0 +41.2 +40.5 +37.1 +18.0 +27.7 +29.4 +11.2 +18.5 +37.5
B-7B-LoRA +20.3 0+7.5 +40.7 +20.7 +21.9 +17.5 +15.9 +23.7 +25.3 0+9.8
B-7B-fft +27.3 +14.8 +37.8 +28.9 +43.3 +13.1 +15.3 +38.5 +34.7 +19.5
@@
Training Strategy
In this study, we follow the two-stage approach of Xu et al. (2023). Unlike traditional DocMT methods, which typically start with parallel sentence training, we explore the effectiveness of this conventional training strategy on LLM-based DocMT models. In this section, we introduce a three-stage training strategy, involving: (1) monolingual document fine-tuning, (2) parallel sentence fine-tuning, and (3) parallel document fine-tuning, for all parameters of the Llama2-7B. The results in Table 5 indicate that the three-stage training strategy is unnecessary for both high-performing languages (Dutch and Romanian) and low-performing languages (Arabic and Chinese) with LLM-based DocMT models.
Scaling Law of Parallel Documents
In this section, we explore the scaling law for fine-tuning parallel documents. We focus on English to German, Romanian, and Chinese translations due to our models’ proficiency. Results for English-German translation are presented in Figure 4, and for English-Romanian and English-Chinese in Appendix G. While LLMs typically excel with minimal training data, different fine-tuning strategies show distinct scaling behaviors. Our LoRA models match full training set performance with just $10\%$ of the data (around $20K$ examples), while fully fine-tuned models achieve near-equivalent performance with only about $1\%$ of the data (approximately $2K$ examples). These insights are crucial for low-resource languages, as recent LLMs are predominantly pre-trained on English text.
Evaluation on Recent Test Sets
Given their pre-training on extensive text corpora, LLMs may be susceptible to data leakage risks. We evaluate our models using recent test sets from WMT2023 (Koehn et al., 2023). These tests, conducted between English and German, not only evaluate the out-of-domain generalization of our models but also help mitigate the risks associated with data leakage. We use spaCy to segment documents and and discard any parallel documents where the source and target sides have a differing number of sentences. Our findings, presented in Table 7, reveal that while Doc2Doc-mT5 models outperform LLM-based models in Table 2, LLM-based models excel in translating out-of-domain text on the WMT2023 test sets. These findings highlight the ability of LLM-based DocMT to generalize well to out-of-domain translation tasks.
@@
In this section, we explore the transferability of translation capabilities acquired from one language pair to others. We assess our English-German LLM-based DocMT models on English-to-other-language test sets, comparing their COMET scores to their base models in Table 6. Our results indicate that models with fine-tuned instructions (Llama2-7B and Bloom-7B) consistently exhibit positive transfer effects across all language pairs, while those with instruction-tuned backbones (Vicuna-7B) benefits only a few languages. These findings suggest that LLMs are more likely to activate their inherent translation abilities during fine-tuning rather than developing new ones.

### v3

@@ Abstract
Large language models (LLMs) have significantly advanced various natural language processing (NLP) tasks. Recent research indicates that moderately-sized LLMs often outperform larger ones after task-specific fine-tuning. This study focuses on adapting LLMs for document-level machine translation (DocMT) for specific language pairs. We first investigate the impact of prompt strategies on translation performance and then conduct extensive experiments using two fine-tuning methods, three LLM backbones, and 18 translation tasks across nine language pairs. Our results show that specialized models can sometimes surpass GPT-4 in translation performance but still face issues like off-target translation due to error propagation in decoding. We provide an in-depth analysis of these LLMs tailored for DocMT, examining translation errors, discourse phenomena, training strategies, the scaling law of parallel documents, recent test set evaluations, and zero-shot crosslingual transfer. Our findings highlight the strengths and limitations of LLM-based DocMT models and provide a foundation for future research.
@@ 1 Introduction
In this study, we analyze moderately-sized LLMs (with $7B$ parameters) across 18 translation tasks involving nine language pairs. We fine-tune three LLMs using Parameter-Efficient Fine-Tuning (PEFT) and Fully Fine-Tuning (FFT). Comparisons with state-of-the-art translation models, using metrics like $s$BLEU, $d$BLEU, and COMET, confirm the superior translation capabilities of LLMs after fine-tuning. However, we identify a significant issue of off-target translations, observed even after exclusive fine-tuning on bilingual corpora. Our further analysis shows that the high off-target translation rate results from the error propagation in the decoding process. Additionally, we present an in-depth analysis of our LLM-based DocNMT models from various perspectives: translation error distribution, discourse phenomena, training strategy, the scaling law of parallel documents, additional evaluations on WMT2023 test sets (Koehn et al., 2023), and zero-shot cross-lingual transfer, aiming to enhance understanding and efficacy of LLMs in DocMT tasks.
@@
  • Selective Excellence in Translation Tasks: Our findings show that our moderately-sized LLMs outperform gpt-4-turbo in certain translation tasks, but struggle in others due to the off-target translation issue due to error propagation in decoding. Despite this, our DocMT models exhibit better context awareness and fewer errors, while maintaining comparable performance.
@@ 2 Related Work
Large language models (LLMs) have demonstrated remarkable proficiency across a wide range of Natural Language Processing (NLP) tasks (Brown et al., 2020; Chowdhery et al., 2022; Scao et al., 2022; Anil et al., 2023; Touvron et al., 2023a, b). Furthermore, recent research has shown that supervised fine-tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF) can significantly enhance their performance when following general language instructions Mishra et al. (2022); Wang et al. (2022); Shen et al. (2023); Li et al. (2023); Wu and Aji (2023); Wu et al. (2024a). More recently, there is a growing body of work exploring the translation capabilities of LLMs (Lu et al., 2023; Zhang et al., 2023; Xu et al., 2023; Robinson et al., 2023; Wu et al., 2024d). However, it is important to note that these efforts have primarily focused on sentence-level machine translation (SenMT) and have not delved into document-level machine translation (DocMT). A noteworthy study in DocMT is conducted by Wang et al. (2023b), where they investigate the document-level translation capabilities of gpt-3.5-turbo, making it the most closely related work to our work.
@@ 3.1 Two-Stage Training
DocMT approaches typically start by pre-training the translation model on parallel sentences and then fine-tuning it on parallel documents (Voita et al., 2019; Ma et al., 2020; Wu et al., 2023). Recently, Xu et al. (2023) proposed a two-stage training strategy. This involves initially fine-tuning a LLM on monolingual text, followed by a second fine-tuning phase on parallel text. Following Xu et al. (2023), our approach begins with fine-tuning an LLM on monolingual documents, followed by fine-tuning on parallel documents.
@@ 3.3 Models
  • State-of-the-art LLMs: For our baseline LLMs in the context of DocMT, we utilize gpt-3.5-turbo and gpt-4-turbo. [^2] Model signatures: gpt-3.5-turbo-1106 and gpt-4-1106-preview. We use the Prompt 4 as detailed in 1(d) during the translation process.
  • Our re-implemented DocMT models: We conduct full fine-tuning on the concatenation-based DocMT model (Tiedemann and Scherrer, 2017), as well as several recent DocMT baselines (Sun et al., 2022; Wu et al., 2023, 2024b), initialized with mT5 (Xue et al., 2021). These models are available with parameters of 300M, 580M, and 1.2B, representing the strong DocMT baseline.
@@ 3.4 Evaluation
We use beam search with the beam size of 5 during translation. As shown in 1(d), previous translations serve as the context for the current translation, so the test examples are translated in their original order, beginning with the first sentence free from context.
@@ 4 A Preliminary Study on Prompts
Table 2: Overall performance on IWSLT2017. # of param. indicates the number of parameters of the model. # of train. param. indicates the number of trainable parameters of the model. All the LLM approaches use Prompt 4 (1(d)) during inference. Best results are highlighted in bold. Best results in each group are underlined.
@@ Results
Our investigation analyzes prompt variations using three PEFT models (L-7B-LoRA, B-7B-LoRA, and V-7B-LoRA) on four English-centric translation tasks involving German and Chinese. Overall results are presented in Table 1. Comparing Prompt 1 (1(a)) and Prompt 2 (1(b)), we find that models fine-tuned with Prompt 2 generally outperform those with Prompt 1, indicating Prompt 2’s effectiveness in enhancing LLM performance. Regarding our second research question (1(a) vs. 1(c)), we observe varied performance. L-7B-LoRA and B-7B-LoRA perform better with Prompt 3, while V-7B-LoRA performs better with Prompt 1. These results highlight varying impacts of prompt variations across models and suggest natural language instructions are less effective when using instruction-tuned language models as model backbones. Finally, LLMs with Prompt 4 (1(d)) achieve the best overall performance, suggesting a positive compound effect of context structure and instructions.
@@ Conclusion
As expected, the prompt plays a significant role in LLM performance. A well-structured prompt, combining appropriate context and clear instructions, can significantly boost model performance. In this work, we use Prompt 4 (1(d)) in our experiments.
@@ 5 Main Results
In our results presented in Table 2, we observe that gpt-4-turbo and gpt-3.5-turbo significantly outperform all other models. Notably, the NLLB variants also demonstrate superior performance among specialized machine translation (MT) models. In the context of DocMT, conventional DocMT models still outperform our LLM-based DocMT models for translations from English to other languages when evaluated using standard MT metrics. Conversely, for translations from other languages to English, our LLM-based DocMT models perform on par or better than conventional DocMT models in $\mu_{s\textsc{BLEU}}$ and $\mu_{d\textsc{BLEU}}$ metrics, while conventional DocMT models maintain superior performance in $\mu_{\textsc{COMET}}$.
@@
As shown in Table 2, models using LoRA generally outperform fully fine-tuned (FFT) LLMs. However, V-7B-fft surpasses V-7B-LoRA in translating other languages to English, likely due to overfitting. Extensive fine-tuning with a large corpus often leads to rapid overfitting, whereas LoRA, which updates only a few parameters, helps prevent this issue. Additionally, L-7B and V-7B models perform similarly, indicating that further fine-tuning with instruction-tuned models does not always improve task-specific performance.
@@
We present the results for translation tasks from other languages to English in Figure 2. For clarity, only the results from our models using LoRA are shown. Our LLM-based DocMT models generally perform better, sometimes even surpassing gpt-4-turbo in certain tasks, though they fail completely in others. A manual review indicates that the primary cause of these failures is off-target translation. An in-depth analysis of this issue is provided in Section 6, and a complete breakdown of the results is available in Appendix E.
6 Analysis
In this section, we examine off-target issues with decoding strategy and use GPT-4 to analyze translation errors. We also explore discourse phenomena, training strategies, the scaling law of parallel documents, and additional evaluations on recent WMT2023 test sets. Additionally, we evaluate the zero-shot crosslingual transfer capabilities of our models and present the results in Appendix I.
@@ 6 Analysis
Table 3: Off-target rate (%) provided by our LLM-based DocMT models for translation tasks from selective languages to English. $\mu_{\%}$ indicates the average off-target rate across all nine language pairs. A lower off-target rate indicates better performance. The complete results are presented in Appendix F.
@@
In Figure 2, our LLM-based DocMT models excel in some translation tasks but struggle in others due to off-target translation issues. We investigate this problem using the fasttext library (Bojanowski et al., 2017) to identify translation languages and quantify off-target rates, which represent the proportion of translations that are off-target. Results are presented in Table 3, with off-target rates reaching up to 98.3% in failing tasks. Notably, only B-7B-LoRA consistently maintains low off-target rates, likely due to Bloom-7B’s multilingual pre-training. Detailed off-target rates are provided in Appendix F.
$s$BLEU $d$BLEU COMET
L-7B-LoRA + ReUse 03.9 04.1 53.9
L-7B-LoRA + ReGen 17.5 19.6 69.7
L-7B-fft + ReUse 02.5 02.6 51.6
L-7B-fft + ReGen 15.9 17.4 65.0
Table 4: Results on Arabic-English translation given by different decoding strategies.
Decoding Strategy
As described in Section 3.4, previous translations provide context for the current translation. We hypothesize that the high off-target translation rate in Table 3 is due to error propagation during the decoding stage. Therefore, we refer to the decoding strategy in Section 3.4 as ReUse, and introduce an alternative strategy where all context translations are re-generated, called ReGen. As shown in Table 4, the ReGen strategy significantly improves translation quality, confirming our hypothesis on decoding error propagation. However, the inference cost of ReGen is four times higher than that of ReUse. These findings highlight the main reason for translation failures in LLM-based DocMT models, offering insights for future research.
002002002002004004004004006006006006008008008008001,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 8002,00020002{,}0002 , 0002,20022002{,}2002 , 2002,40024002{,}4002 , 400MistranslationOvertranslationUndertranslationAdditionOmissionGrammarCohesionCoherenceInconsistent styleMultiple terms in translation1,39513951{,}3951 , 3957177177177171,22412241{,}2241 , 2248408408408404844844844842342342342347097097097094544544544547257257257253463463463461,35613561{,}3561 , 3567177177177171,35813581{,}3581 , 3588638638638636166166166162662662662667277277277275655655655657717717717713393393393391,71217121{,}7121 , 7127157157157159779779779772452452452454174174174174864864864868048048048045925925925928248248248242732732732732,00220022{,}0022 , 0028368368368361,55115511{,}5511 , 551586586586586868868868868474474474474933933933933717717717717854854854854386386386386# of occurencesL-7B-LoRAL-7B-fftGoogleTransDoc2Doc-mT5-1.2B
Figure 3: Error type analysis given by gpt-4-turbo for translations from English to German, Romanian, and Chinese. The error types in orange are context-dependent. We omit those error types that are rare.
@@
To understand the translation capabilities of our LLM-based DocMT models, we select specific error types from the Multidimensional Quality Metrics (MQM) framework (Burchardt, 2013). Kocmi and Federmann (2023) demonstrate that gpt-4 can identify error spans and achieve state-of-the-art MT evaluation accuracy, so we use gpt-4-turbo to analyze translation errors in texts translated by these models. Due to resource constraints, we focus on four models: L-7B-LoRA, L-7B-fft, Doc2Doc-mT5-1.2B, and GoogleTrans, assessing translations from English to German, Romanian, and Chinese. The error identification prompt is detailed in Appendix D, and we present the frequency of error types in Figure 3. Most errors are limited to individual sentences. Despite similar scores in metrics such as $s$BLEU, $d$BLEU, and COMET among the models, our LLM-based DocMT models (L-7B-LoRA and L-7B-fft) exhibit fewer context-independent and context-dependent errors. This suggests that current evaluation metrics may not sufficiently assess document-level translations and indicates that fine-tuning LLMs holds promise for enhancing DocMT performance.
En-De En-Fr
Doc2Doc-mT5-1.2B 77.0 89.9
L-7B-LoRA 83.1 95.1
L-7B-fft 81.1 94.6
B-7B-LoRA 75.5 91.9
B-7B-fft 68.3 90.8
V-7B-LoRA 84.9 94.8
V-7B-fft 84.4 94.8
Table 5: Accuracy (in %) on the English-German and English-French contrastive test sets. Best results are highlighted in bold.
@@
To evaluate our LLM-based DocMT model’s ability to leverage contextual information, we assessed it using the English-German contrastive test set by Müller et al. (2018) and the English-French contrastive test set by Lopes et al. (2020). This evaluation tests the model’s accuracy in selecting the correct pronoun from multiple translation options. Results, shown in Table 5, reveal that models initialized with Llama2-7B and Vicuna-7B outperform Doc2Doc-mT5-1.2B, while Bloom-7B-initialized models perform worse, indicating that contextual understanding is mostly acquired during pre-training, as detailed by Scao et al. (2022) due to the lack of German text in Bloom pre-training. The generative accuracy results following Post and Junczys-Dowmunt (2023) are presented in Appendix G.
@@
Table 6: Results from one-stage, two-stage, and three-stage training strategies. The one-stage strategy involves directly fine-tuning Llama2-7B on parallel documents. The two-stage results are produced by L-7B-fft. In the three-stage strategy, all model parameters of Llama2-7B are fine-tuned across all three stages.
Training Strategy
In this study, we follow the two-stage approach of Xu et al. (2023). Unlike traditional DocMT methods that typically start with parallel sentence training, we investigate the effectiveness of this strategy on LLM-based DocMT models. Specifically, we attempt to directly fine-tune the LLMs on parallel documents (one-stage) and add an extra fine-tuning stage using parallel sentences to the training strategy in Section 3.1 (three-stage). The results in Table 6 indicate that both the one-stage and three-stage training strategies are sub-optimal for both high-performing languages (Dutch and Romanian) and low-performing languages (Arabic and Chinese) with LLM-based DocMT models.
1102030405060708090100404040406060606080808080Percentage (%) of training dataCOMETL-7B-LoRAL-7B-fftB-7B-LoRAB-7B-fftV-7B-LoRAV-7B-fft
@@
Scaling Law of Parallel Documents
We explore the scaling law for fine-tuning parallel documents. Results for English-German translation are presented in Figure 4, and for English-Romanian and English-Chinese in Appendix H. While LLMs typically excel with minimal training data, different fine-tuning strategies show distinct scaling behaviors. Our LoRA models match full training set performance with just $10\%$ of the data (around $20K$ examples), while fully fine-tuned models achieve near-equivalent performance with only about $1\%$ of the data (approximately $2K$ examples). These are crucial for low-resource languages, as recent LLMs are mainly pre-trained on English text.
Evaluation on Recent Test Sets
Given their extensive pre-training on large text corpora, LLMs are susceptible to data leakage risks. We evaluate our models using recent test sets from WMT2023 (Koehn et al., 2023). These tests, conducted between English and German, assess the out-of-domain generalization of our models and help mitigate data leakage risks. We use spaCy to segment documents and discard any parallel documents where the source and target sides have a differing number of sentences. Our findings, presented in Table 7, reveal that while DocNMT5 models outperform LLM-based models in Table 2, LLM-based models excel in translating out-of-domain text on the WMT2023 test sets. These results highlight the ability of LLM-based DocNMT to generalize well to out-of-domain translation tasks.
@@
$\mu_{\Delta}$ Ar De Fr Zh
L-7B-LoRA +29.4 +36.3 +38.8 +37.2 +29.5
L-7B-fft +29.0 +41.2 +40.5 +37.1 +37.5
B-7B-LoRA +20.3 0+7.5 +40.7 +20.7 +9.8
B-7B-fft +27.3 +14.8 +37.8 +28.9 +19.5
V-7B-LoRA 0-8.9 -12.6 +22.1 +18.9 -34.1
V-7B-fft 0-1.4 0+7.3 +25.2 +17.7 0-3.5
Table 8: The difference ($\Delta$) in COMET scores on the test sets from English to other languages between our English-German LLM-based DocMT models and their backbones. $\mu_{\Delta}$ indicates the average difference across all the languages in this table. The complete results are presented in Appendix I.
@@
We also explore the transferability of translation capabilities acquired from one language pair to others. We assess our English-German LLM-based DocMT models on English-to-other-language test sets, comparing their COMET scores to their base models in Table 8. Our results indicate that models with fine-tuned instructions (Llama2-7B and Bloom-7B) consistently exhibit positive transfer effects across all language pairs, while those with instruction-tuned models (Vicuna-7B) benefits only few languages.